previous 5 Part A What do you spectly for alternators magnetic field? HNT Notice that rotation rate ofthe coil is given in revs Perminutes, not revs per Youre an electrical engineer designing an aaemator (the generator that charges a cars banery) Mechanical engineers a cm diameter rotating cod, and you detemine that you can fit 250 tumsin coil To charge a Express your answer to twosignincantsgures and inelude the appropriate unis. battery, you need a peak output of 16 V when the atemator B. Value Units Submit My Answers Esetes Continue mmmm(mmmmmmmmmmmmmmmmmmmmmmmmmmmmmmmmmm
Answer #1

Vmax = 16 V, N = 250 , d =,10 cm , r = 5 cm

w = 1100 rpm = 1100*2*3.14/60 = 115.13 rad/s

incuded emf Vmax = NABw

16 = 250*3.14*0.05^2*115.13*B

B = 0.071 T